FOURIER TRANSFORM EMISSION SPECTROSCOPY OF THE G ^{3}Φ − X ^{3}Φ, C ^{3}Δ − X ^{3}Φ AND G ^{3}Φ − C ^{3}Δ SYSTEMS OF CoCl

\underline{R. S. RAM}, Department of Chemistry, University of Arizona, Tucson, AZ 85721; I. GORDON, Department of Physics, University of Waterloo, Waterloo, Ont., Canada N2L 3G1; T. HIRAO, S. YU, P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ont., Canada N2L 3G1; and B. PINCHEMEL, Laboratoire PhLAM, UMR CNRS 8523, Centre d’Etudes et de Recherches Lasers et Applications, Université des Sciences et Technologies de Lille, 59655 Villeneuve d’Ascq Cedex, France.

The emission spectrum of CoCl has been recorded in the 2000–20000 cm^{−1} region using a Fourier transform spectrometer. The bands were excited in a carbon tube furnace by heating CoCl_{2} to a temperature of about 2200^{∘}C. The bands observed in the 2000–12000 cm^{−1} interval have been classified into three transitions: G ^{3}Φ−X ^{3}Φ, C ^{3}Δ−X ^{3}Φ and G ^{3}Φ−C ^{3}Δ, analogous to the near infrared transitions of CoF [R. S. Ram, P. F. Bernath and S. P. Davis, J. Chem. Phys. 104, 6949 (1996)]. Two sub-bands with 0–0 origins near 10486.9 and 10449.9 cm^{−1} have been assigned as the ^{3}Φ_{3} −^{3} Φ_{3} and 3Φ_{4} −^{3} Φ_{4} sub-bands of the G ^{3}Φ − X ^{3}Φ transition. Two other 0–0 bands with origins near 2636.6 and 2726.7 cm^{−1} have been assigned as ^{3}Δ_{2} −^{3} Φ_{3} and ^{3}Δ_{3} −^{3} Φ_{4} sub-bands of the C ^{3}Δ − X ^{3}Φ transition. In addition, two very weak bands observed with origins near 7850.3 and 7723.2 cm^{−1} have been identified as the 0–0 bands of the ^{3}Φ_{3} −^{3} Δ_{2} and ^{3}Φ_{4} −^{3} Δ_{3} sub-bands of the G ^{3}Φ − C ^{3}Δ transition. A rotational analysis of a number of vibrational bands of these transitions has been obtained and spectroscopic constants for the low-lying electronic states of CoCl have been determined.